OpenStudy (anonymous):

Garrett throws a dart at a circular dart board. The dart board has a radius of 15 inches, and the bull’s eye in the center of the dart board has a radius of 2 inches. What is the probability that a dart thrown at random within the dartboard will hit the bull’s eye? Round your answer to the nearest tenth, if necessary.

OpenStudy (whpalmer4):

Area of a circle with radius \(r\): \(A = \pi r^2\) Find the area of the bull's eye. Find the area of the dart board. The ratio of the areas is the probability that a randomly placed dart will hit the bull's eye.

OpenStudy (anonymous):

The answer choices are in % though. A 1.8% B 5.6% C 8.4% D 13.3%

OpenStudy (whpalmer4):

Okay, multiply the probability (which should be somewhere between 0 and 1) by 100 and tack on a percent sign...

OpenStudy (anonymous):

I'm confused. So 15^2pi = about 706.9 and 2^2pi = about 12.6

OpenStudy (anonymous):

I know the answer isn't A.

OpenStudy (anonymous):

@tcarroll010 Do you know how to solve this?

OpenStudy (anonymous):

The P(hitting bull's eye) = (A)rea of bull's eye divided by (A)rea of dartboard P = [(pi)(2^2)] / [(pi)(15^2)] = 4 / 225 = (4 / 225) (100) % Just do the division and you'll get your answer.

OpenStudy (anonymous):

All good now, @savannahp8 ?

OpenStudy (anonymous):

Wouldn't you have to multiply by pi? So (706.9/12.6) (100) % = 5.6% My friend got 1.8% and got it wrong so it must be 5.6% Right?

OpenStudy (anonymous):

pi is in the numerator and the denominator, so it cancels. P = (4 / 225) (100) % = 1.77777777 % approx. 1.8 % You're done!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @savannahp8

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Yep! Thank you.

OpenStudy (anonymous):

uw! Have a good day!

OpenStudy (anonymous):

You too!!